# Effect of realistic oscillator phase noise on the performance of cell-free networks

### v1

@@ II-B4 Received Signal Model
where $X_{k,i}\in\mathbb{C}$ from UE $k$ with power $p_{k}$, $\mathbf{y}_{i}=[\mathbf{y}_{1,n}^{T}\cdots\mathbf{y}_{M,n}^{T}]^{T}\in\mathbb{C}^{W\times 1}$ is a block vector, while $\mathbf{h}_{k}=[\mathbf{h}_{k1}^{T}\cdots\mathbf{h}_{kM}^{T}]^{T}\sim\mathcal{CN}(\mathbf{0},\mathbf{R}_{k})$ is the concatenated channel vector from all APs and $\mathbf{\Theta}_{k}=\mathrm{diag}(\mathbf{\Theta}_{1k},\cdots,\mathbf{\Theta}_{Mk})\in\mathbb{C}^{W\times W}$ is the phase noise matrix including CPE for this OFDM symbol. The block diagonal spatial correlation matrix $\mathbf{R}_{k}=\mathrm{diag}(\mathbf{R}_{k1},\cdots,\mathbf{R}_{kM})\in\mathbb{C}^{W\times W}$ is obtained by assuming that the channel vectors of different APs are independently distributed. Finally, the thermal noise vector $\mathbf{n}\sim\mathcal{CN}(0,\sigma^{2}_{W}\mathbf{I}_{W})$ is spatially and temporally independent and has variance $\sigma_{W}$. We define the effective channel for the ${\mathfrak{t}}$-th OFDM symbol (including the CPE) as $\tilde{\mathbf{h}}^{\mathfrak{t}}=\mathbf{\Theta}^{\mathfrak{t}}\mathbf{h}$.
As we stated before, although all APs receive the signals from all UEs, only a subset of the APs contributes to signal detection. Thus, the network estimates of $X_{k,i}$ are
@@
In this paper, we use Maximum Ratio (MR) and Minimum Mean-Square Error (MMSE) given as
@@
III Phase Noise Model
In this section, we will first present a simplified PN phase model commonly found in the literature (Sec. III-A). Then, to create a realistic PN model we refer to the practical implementation of the Software Defined Radio (SDR) transceivers. In Sec. III-C, we briefly describe the contributions to PN introduced by different components of an LO. We conclude the section with presenting data of the real devices we use for the modeling in Sec. III-D.
@@ III-A Simplified PN Model: Free Running Oscillator
As a reference, we use a model where the phase follows Weiner process [10] described by $\phi_{m}(t)\sim\mathcal{N}(\phi_{m}(t-1),\sigma^{2}_{\phi})$ and $\varphi_{k}(t)\sim\mathcal{N}(\varphi_{k}(t-1),\sigma^{2}_{\varphi})$, where $t$ denotes the channel use number while the variance of innovation can be expressed by [10]
@@ III-B Hardware-Inspired Phase Noise Model
To make the PN modeling more accurate and realistic, we implement a model of real-world devices. We consider SDRs transceivers which use LOs used for frequency up-/downconversion and sample I/Q at baseband. In our model, a LO is a voltage-controlled oscillator (VCO) which is locked to a reference crystal oscillator (XO) by a phase-locked loop (PLL). This model allows us to precisely model the behavior of many common SDRs which are based on integrated transceiver chips (e.g., Analog Devices AD9361). We also apply this model to more complex SDRs which may have multiple PLL stages to filter the reference clock and generate the sampling and carrier frequency (e.g. Ettus USRP X300 series). In the latter case, we limit our consideration to the last VCO+PLL (the one that generates the actual carrier frequency) and to one XO.
III-C Phase Noise Model of a Local Oscillator
To describe a PN of a standalone oscillator, the engineering community uses one-sided Power Spectral Density (PSD) $S(\delta f)$ where $\delta f$ is the frequency offset from the carrier frequency. Depending on the considered time scale, $\delta f$ can go down to fractions of Hertz (corresponding to time intervals of a second and more) and up to hundredss of MHz (when we consider period-to-period jitter of a carrier).
Different sources of PN prevail at different time scales: white and flicker noises of phase at high $\delta f$, white and flicker noises of frequency, and frequency random walk at slowest time scales. An example of contributions of these noise types can be seen in Fig. 3.
Naturally, XOs have good stability at long time scales and moderate PN at high $\delta f$. At the same time, they cannot be used for directly generating frequencies over $\sim 100$ MHz. The frequency can be multiplied to get the carrier frequency $f_{c}$, but the PN amplitude increases proportionally to the carrier frequency as
@@ III-C Phase Noise Model of a Local Oscillator
On the other side, VCOs can generate high-frequencies directly and have low phase noises at high $\delta f$, but their frequency is very unstable.
By using PLL, it is possible to obtain a high-frequency oscillator that has i) low PN of VCO at high-frequency offsets and ii) low PN of XO at low-frequency offsets (ensuring good long- and short-term stability). At Fig. 4, a block diagram of a PLL is shown. The phase detector takes two inputs: a reference frequency and a divided VCO frequency, and compares the phase difference. The phase difference signal is then low-pass filtered and used to tune the VCO frequency.
@@
The low-pass filter removes the high-frequency noise component of the XO. Thus, at the frequency offsets higher than the cut-off frequency $f_{\mathrm{PLL}}$ the PN is defined by the VCO. At the frequency offsets below $f_{\mathrm{PLL}}$ the LO PN is defined by the PN of the XO (amplified according to (15)) and the PN contribution of the phase detector. The total PN of the LO is
@@
In our simulation, we take closed-loop performance data of PLL+VCO $S_{\mathrm{PLL+VCO}}(\delta f)$ from the datasheets of the particular integrated circuits (ICs), and use it as the input parameters for our model. This data combines the contributions of the VCO noise and phase detector noise.
III-D Real Devices and Their Performance
In this work, we are interested in oscillators’ stabilities at time scales of the 5G NR protocol: from the duration of an OFDM symbol $T_{\mathrm{OFDM}}=71.4$ $\mu$s to the duration of a frame $T_{\mathrm{c}}=1$ ms. Thus, our region of interest is
With (16), we compute the LO phase noises at these frequency offsets for the following devices. The resulting PN figures are given in Table I.
III-D1 NI USRP B200
As an example of a simple SDR, we selected NI USRP B200. It has a local thermal-compensated XO (TCXO) reference which is fed to Analog Devices AD9361 transceiver IC including its own PLL+VCO.
III-D2 NI USRP RIO 2954R
As an example of an advanced SDR, we consider the USRP RIO 2954R which is an SDR based on Ettus X300 series, which includes a clock cleaning and distribution IC with two additional PLL loop stages: one with low-noise voltage controlled XO (LN-VCXO) and the second one with high-frequency low-noise VCO. The latter frequncy is then divided (if required) and used for clocking ADC/DAC and as a reference for Maxim Integrated MAX2871 LO.
To simplify our consideration, we use the free-running performance of the LN-VCXO as the reference XO, and the PN of the MAX2871.
III-D3 Instrumental Grade Transceiver
As a reference, we consider the performance of Texas Instruments LMX2594 high-performance RF synthesizer IC locked to Symmetricom 1000C low noise and high short-term stability oven-controlled XO (OCXO).
@@ III-D3 Instrumental Grade Transceiver
TABLE I: Calculated phase noise of the considered devices at different frequncy offsets. Values at offsets from 10 Hz to 100 kHz are used for simulating the cell-free scenario with the considered 5G NR numerology.
@@ IV Numerical results
Aligned with selection of FR1, we consider $M=100$ APs with $L=4$ antennas and $K=40$ UEs in an $0.4\times 0.4$ km2 area. Large-scale propagation effects are modeled as
$\beta_{mk}(d)=\Lambda_{0}+10\eta\log(d/d_{0})+\Lambda_{sh},$ (18)
where $d$ is the distance between AP $m$ and UE $k$, $\Lambda_{0}=-35.3$ is the PL at reference distance $d_{0}=1$ m, pathloss exponent $\eta=3.76$, and $\Lambda_{sh}$ is lognormally distributed with the standard deviation of $\sigma_{sf}=10$. Moreover, UEs transmit with the same power $p=p_{k}=p_{m}=100$ mW in both uplink training and transmission phases. The thermal noise variance is $\sigma_{N}$= -174 dBm/Hz. Regarding the PN model, we use the same parameters as in [7, 8] and (14) for our time scales of one OFDM symbol $T_{\rm OFDM}$, the obtained variance of the innovation equals $0.23$ (note that [9] used 4 times higher variance resulting in even more severe PN).
@@
Fig. 5 illustrates the difference between PN produced by the model used in [7, 8] and in our study. It is obvious that realistic LOs are substantially more stable than the Free Running Oscillator models used in prior art on CF. Realistic LOs drift moderately resulting in very small CPE while the previously used model produced PN causing significant CPE and ICI effects [9].
@@
Fig. 6 shows the variation of the achievable SE per user over OFDM symbols due to the evolving CPE. Under severe PN generated by models parameterized as in [7, 8], the achievable SE decreases within one TTI since the aggregate contribution from PN becomes higher. However, when the realistic PN model is used (we use the model of NI USRP B200 as it has the least stable LO), we observe no SE degradation within the TTI (1 ms).
@@ V Conclusions
This study critically examined the performance of Cell-Free (CF) networks under the influence of Phase Noise (PN) using a realistic hardware-based model. Unlike previous works that employed simplified oscillator models, our approach leverages a more accurate representation of Local Oscillators (LO) used in radios like NI USRP B200 and NI USRP RIO. Our findings reveal that even less stable, cost-effective LOs maintain sufficient phase stability to prevent Spectral Efficiency (SE) losses over the standardized 3GPP Transmission Time Interval (TTI) of 1 ms. This substantiates the viability of employing existing 5G standards in the deployment of scalable CF networks for future 6G infrastructure. Furthermore, the resilience of CF networks to hardware imperfections, as demonstrated by our simulations, supports the practical implementation of these networks, potentially alleviating some of the concerns raised by previous theoretical studies.

### v2

@@ II-B4 The Received Signal Model
where $X_{k,i}\in\mathbb{C}$ represents the signal from UE $k$ with power $p_{k}$, and $\mathbf{y}_{i}=[\mathbf{y}_{1,n}^{T}\cdots\mathbf{y}_{M,n}^{T}]^{T}\in\mathbb{C}^{W\times 1}$ is a block vector. The concatenated channel vector from all APs is denoted as $\mathbf{h}_{k}=[\mathbf{h}_{k1}^{T}\cdots\mathbf{h}_{kM}^{T}]^{T}\sim\mathcal{CN}(\mathbf{0},\mathbf{R}_{k})$ is the phase noise matrix, including the CPE for this OFDM symbol.
The block diagonal spatial correlation matrix $\mathbf{R}_{k}=\mathrm{diag}(\mathbf{R}_{k1},\cdots,\mathbf{R}_{kM})\in\mathbb{C}^{W\times W}$ is obtained under the assumption that the channel vectors from different APs are independently distributed. Finally, the thermal noise vector $\mathbf{n}\sim\mathcal{CN}(0,\sigma^{2}_{W}\mathbf{I}_{W})$ is spatially and temporally independent, with a variance of $\sigma_{W}$. We define the effective channel for the $t$-th OFDM symbol (including the CPE) as $\tilde{\mathbf{h}}^{\mathfrak{t}}=\mathbf{\Theta}^{\mathfrak{t}}\mathbf{h}$.
As we stated before, although all APs receive the signals from all UEs, only a subset of the APs contributes to signal detection. Thus, the network estimates of the signal $X_{k,i}$ transmitted by user $k$ over subcarrier $i$ can be written as
@@
In this paper, we use MR and MMSE given as
@@
III The Phase Noise Model
In this section, we begin by presenting a simplified PN model that is commonly referenced in the literature (Sec. III-A). Next, we develop a more realistic PN model based on the practical implementation of Software Defined Radio (SDR) transceivers. In Sec. III-C, we briefly discuss the contributions to PN introduced by various components of a LO. We conclude this section by presenting data from the actual devices we utilize for modeling in Sec. III-D.
@@ III-A Simplified PN Model: Free Running Oscillator
As a reference, we employ a model in which the phase follows a Wiener process [10], described by $\phi_{m}(t)\sim\mathcal{N}(\phi_{m}(t-1),\sigma^{2}_{\phi})$ and $\varphi_{k}(t)\sim\mathcal{N}(\varphi_{k}(t-1),\sigma^{2}_{\varphi})$, where $t$ indicates the channel use number, while the variance of innovation is given by
@@ III-B Hardware-Inspired Phase Noise Model
For accurate and realistic PN modeling, we implement a model of real-world devices. We consider SDR transceivers which use LOs for frequency up-/down conversion and sample I/Q at baseband. In our model, a LO is a voltage-controlled oscillator (VCO) which is locked to a reference crystal oscillator (XO) by a phase-locked loop (PLL).
This model allows us to precisely model the behavior of many common SDRs which are based on integrated transceiver chips (e.g., Analog Devices AD9361). We also apply this model to more complex SDRs which may have multiple PLL stages to filter the reference clock and generate the sampling and carrier frequency (e.g. Ettus USRP X300 series). In the latter case, we limit our consideration to the last VCO+PLL (the one that generates the actual carrier frequency) and to one XO.
III-C The Phase Noise Model of a Local Oscillator
To describe the PN of a standalone oscillator, the engineering community employs one-sided Power Spectral Density (PSD) denoted as $S(\delta f)$ where $\delta f$ represents the frequency offset from the carrier frequency. Depending on the time scale under consideration, $\delta f$ can range from fractions of Hertz—corresponding to time intervals of one second or more—to hundreds of MHz, particularly when analyzing period-to-period jitter of a carrier. Various sources of PN dominate at different time scales: at high $\delta f$ white and flicker noise of phase are prevalent, while at slower time scales, white and flicker noise of frequency, as well as frequency random walk, become significant. An illustration of the contributions from these noise types is shown in Fig. 3.
Crystal oscillators exhibit good stability over long time scales and moderate phase noise at high $\delta f$. However, they are not suitable for directly generating frequencies exceeding approximately $\sim 100$ MHz. When the frequency is multiplied to the carrier frequency $f_{c}$, the amplitude of phase noise increases proportionally with the carrier frequency:
@@ III-C The Phase Noise Model of a Local Oscillator
In contrast, VCOs can directly generate high frequencies and exhibit low phase noise at high $\delta f$, but their frequency stability is notably poor. To create a frequency source that combines both good stability and low noise, XOs and VCOs are often integrated using PLL.
By utilizing a PLL, it is possible to obtain a high-frequency oscillator that exhibits: i) low PN from the VCO at high-frequency offsets, and ii) low PN from the XO at low-frequency offsets, ensuring both long- and short-term stability. Fig. 4 illustrates a block diagram of a PLL. The phase detector receives two inputs: a reference frequency and a divided VCO frequency, which it uses to compare the phase difference. The resulting phase difference signal is then low-pass filtered to tune the VCO frequency. This low-pass filter effectively removes the high-frequency noise component from the XO.
@@
Consequently, at frequency offsets greater than the cut-off frequency $f_{\mathrm{PLL}}$, the PN is primarily determined by the VCO. Conversely, at frequency offsets below $f_{\mathrm{PLL}}$, the LO PN is influenced by the PN of the XO (amplified according to equation (15)) and the PN contribution from the phase detector. The total PN:
@@
For our simulations, we gather closed-loop performance data of the PLL combined with the VCO, denoted as $S_{\mathrm{PLL+VCO}}(\delta f)$, from the datasheets of specific integrated circuits (ICs). This data encapsulates the contributions from both VCO noise and phase detector noise.
@@
TABLE I: The calculated phase noise of the considered devices at different frequency offsets. Values at offsets from 10 Hz to 100 kHz are used for simulating the cell-free scenario with the considered 5G NR numerology.
III-D Real Devices and Their Performance
In this work, we focus on the stability of oscillators over time scales relevant to the 5G NR protocol, ranging from the duration of an OFDM symbol $T_{\mathrm{OFDM}}=71.4$ $\mu$s to the duration of a frame $T_{\mathrm{c}}=1$ ms. Therefore, our area of interest is defined by the frequency range
With (16), we compute the LO phase noises at these frequency offsets for the following devices described in Table I.
III-D1 NI USRP B200
We selected the NI USRP B200 as an example of a simple Software-Defined Radio. It features a locally thermal-compensated crystal oscillator (TCXO) reference that feeds into the Analog Devices AD9361 transceiver IC, which includes its own PLL and VCO.
III-D2 NI USRP RIO 2954R
For a more advanced SDR, we consider the USRP RIO 2954R, which is based on the Ettus X300 series. This device incorporates a clock cleaning and distribution IC with two additional PLL loop stages: one utilizing a low-noise voltage-controlled crystal oscillator (LN-VCXO) and the other employing a high-frequency, low-noise VCO. The latter frequency can be divided as needed and is used for clocking the ADC/DAC and as a reference for the Maxim Integrated MAX2871 LO. For simplicity in our model, we utilize the free-running performance of the LN-VCXO as the reference XO and the PN of the MAX2871.
III-D3 The Instrumental Grade Transceiver
As a reference point, we evaluate the performance of the Texas Instruments LMX2594 high-performance RF synthesizer IC, which is locked to the Symmetricom 1000C low-noise, high short-term stability oven-controlled crystal oscillator (OCXO).
@@ IV Numerical Results
In alignment with the selection of FR1, we consider $M=100$ APs equipped with $L=4$ antennas and $K=40$ UEs within $0.4\times 0.4$ km2 area. Large-scale propagation effects are modeled as $\beta_{mk}(d)=\Lambda_{0}+10\eta\log(d/d_{0})+\Lambda_{sh},$ where $d$ is the distance between AP $m$ and UE $k$, $\Lambda_{0}=-35.3$ is the PL at reference distance $d_{0}=1$ m, pathloss exponent $\eta=3.76$, and $\Lambda_{sh}$ is lognormally distributed with the standard deviation of $\sigma_{sf}=10$. Moreover, UEs transmit with the same power $p=p_{k}=p_{m}=100$ mW in both uplink training and transmission phases. The thermal noise variance is $\sigma_{N}$= -174 dBm/Hz. Regarding the PN model, we use the same parameters as in [7, 8] and (14) for our time scales of one OFDM symbol $T_{\rm OFDM}$, the obtained variance of the innovation equals $0.23$ (note that [9] used 4 times higher variance resulting in even more severe PN).
@@
Fig. 5 illustrates the differences in PN produced by the models used in [7, 8] compared to our study. The realistic LOs are significantly more stable than the free-running oscillator models employed in previous research on CF systems. The realistic LOs exhibit moderate drift, resulting in an order of magnitude smaller CPE than the earlier models [7, 8]. Moreover, ICI is negligible when using realistic LOs, whereas previous theoretical models generated PN that caused significant ICI.
@@
Fig. 6 depicts the variation in achievable SE per user across OFDM symbols due to the evolving CPE. Under severe PN conditions, as parameterized in models [7, 8], the achievable SE decreases by by 10% within one TTI due to the cumulative effect of PN drift. However, when employing the realistic PN model (based on the NI USRP B200, which has the least stable LO), we observe no degradation in SE within the TTI (1 ms).
@@ V Conclusions
This paper critically examines the performance of cell-free (CF) massive MIMO (mMIMO) systems under the influence of phase noise (PN) using a realistic hardware-based model. Unlike previous studies that relied on simplified oscillator models, our approach utilizes a more accurate representation of local oscillators (LOs) found in radios such as the NI USRP B200 and NI USRP RIO. While this paper focuses on CF mMIMO systems, the PN model can be applied to evaluate the performance of any RF system operating within the sub-6 GHz frequency range (FR1). The proposed OFDM CF mMIMO model with realistic PN serves as a valuable tool for assessing and designing practical CF systems, such as defining numerology and pilot allocation schemes. Our findings indicate that even less stable, cost-effective LOs maintain adequate phase stability to prevent Spectral Efficiency losses over the standardized 3GPP transmission time interval of 1 ms. These results affirm the feasibility of using existing 5G standards in the deployment of scalable CF mMIMO systems for future 6G infrastructure. The demonstrated resilience of CF mMIMO systems to hardware imperfections supports the practical implementation of these networks.